OpenStudy (anonymous):

Solve x – 10 = –15..Is the answer -5? thank you

OpenStudy (anonymous):

yes you are correct

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

anytime

OpenStudy (anonymous):

yes it is :)

OpenStudy (anonymous):

we want to have x all by itself on the left, so add 10 to the left side. now you have to add 10 to the right side too (have to do the same thing to both sides) x-10+10=-15+10 x+0=-5 x=-5